OpenStudy (anonymous):

1688/300 how do I work this out by hand? no calculator...

terenzreignz (terenzreignz):

\[\Large \frac{1688}{300}\] It's best we go slowly, cancel out common factors one at a time... Could you name a common factor between 1688 and 300?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so divide both by 2

terenzreignz (terenzreignz):

Yup. we get... \[\Large \frac{\cancel{1688}^{844}}{\cancel{300}^{150}}=\frac{844}{150}\] Any more common factors?

OpenStudy (anonymous):

sorry it should of been 300/1688

terenzreignz (terenzreignz):

Doesn't matter, let's just reverse them \[\Large \frac{300}{1688}= \frac{150}{844}\] Any more common factors?

OpenStudy (anonymous):

4

terenzreignz (terenzreignz):

4 is not a factor of 150, is it?

OpenStudy (anonymous):

no it isnt

terenzreignz (terenzreignz):

So not 4. Anything else?

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

My advice at this stage, keep it simple, and stick to primes...

terenzreignz (terenzreignz):

Okay, 2 again. \[\Large \frac{\cancel{150}^{75}}{\cancel{844}^{422}}=\frac{75}{422}\] Any more common factors?

OpenStudy (anonymous):

yep, now we get into decimals is that right?

terenzreignz (terenzreignz):

You're supposed to get it into decimals? D:

OpenStudy (anonymous):

2 again would give 37.5/211

terenzreignz (terenzreignz):

I'd have thought we should stop at 75/422

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

I need to know what it is all the way down to decimal, how do I prove it down to 0.17

OpenStudy (anonymous):

its too hard right.

OpenStudy (anonymous):

thanks I have it sorted, my lecturer said I can manipulate the numbers by rounding up and down, I can square it up by making the 1st one 300/1700 which is easier, could you help from there.

terenzreignz (terenzreignz):

I guess. There's a pretty obvious common factor between these two. (HINT: They both end with a double zero)

OpenStudy (anonymous):

100

OpenStudy (anonymous):

3/17

terenzreignz (terenzreignz):

Those are both prime, but I take it you can handle it from there?

OpenStudy (anonymous):

I guess I could make it 4/16 by what my lecturer is saying??

OpenStudy (anonymous):

or 3/15

OpenStudy (anonymous):

1/3 0.33

terenzreignz (terenzreignz):

I don't know what you're allowed/supposed to do with it. As far as I know, 3/17 is simplified enough

OpenStudy (anonymous):

yep thanks I can do long division from there. 3/17 is 0.17 I know how to arrive at that answer thankyou.

terenzreignz (terenzreignz):

Okay then. ^_^